OpenStudy (anonymous):

Most car owners drive their cars on an average of 12,000 miles a year. Assuming 20miles/gallon is the gas consumption rate, determine the amount of fuel consumed by 150 million car owners on the following time basis: (Express your results in gallons and liters). a) average daily basis b) average monthly basis c) average yearly basis d) average over 10 years

OpenStudy (anonymous):

12000*20=240,000 gallons consumed for one person a year?

OpenStudy (anonymous):

3.6x10^13 gallons a year consumed by 150 million car owners at an average 12,000 miles/gallon a year of one car owner. 150 million ppl * 12000=1.8*10^12 miles a year by 150 million people 1.8*10^12 miles * 20 mi/gallon= 3.6*10^13 gallons used per year by 150 million car owners? is this correct?

OpenStudy (anonymous):

Wow thats long! Lol.

OpenStudy (anonymous):

For part d. I got 3.6*10^14 gallons consumed by 150 million car drivers in 10 years.